### v3

Abstract
This study presents alternating optimization (AO) algorithms for computing $\alpha$-mutual information ($\alpha$-MI) and $\alpha$-capacity based on variational characterizations of $\alpha$-MI using a reverse channel. Specifically, we derive several variational characterizations of Sibson, Arimoto, Augustin–Csiszár, and Lapidoth–Pfister MI and introduce novel AO algorithms for computing $\alpha$-MI and $\alpha$-capacity; their performances for computing $\alpha$-capacity are also compared. The comparison results show that the AO algorithm based on the Sibson MI’s characterization has the fastest convergence speed.
I Introduction
Sibson mutual information (MI) $I_{\alpha}^{\text{S}}(X;Y)$ [1], Arimoto MI $I_{\alpha}^{\text{A}}(X;Y)$ [2], Augustin–Csiszár MI $I_{\alpha}^{\text{C}}(X;Y)$ [3],[4], and Lapidoth–Pfister MI $I_{\alpha}^{\text{LP}}(X;Y)$ [5],[6] are well-known generalizations of the Shannon MI $I(X;Y)$ [7] using the tunable parameter $\alpha\in(0,1)\cup(1,\infty)$. These types of MI are referred to as $\alpha$-mutual information ($\alpha$-MI) [8] and defined based on the Rényi entropy $H_{\alpha}(X)$, Arimoto conditional entropy $H_{\alpha}^{\text{A}}(X|Y)$, and Rényi divergence $D_{\alpha}(\cdot||\cdot)$. The maximum value of the $\alpha$-MI for the input distribution $p_{X}$, given by $C_{\alpha}^{(\cdot)}:=\max_{p_{X}}I_{\alpha}^{(\cdot)}(X;Y)$, is referred to as $\alpha$-capacity. The $\alpha$-MI and $\alpha$-capacity characterize various problems in information theory, such as the error and correct decoding exponents in channel coding [9, 10, 11, 12, 13, 14], the generalized cutoff rate in channel coding [4], the error and strong converse exponents in composite hypothesis testing [6], and the (maximal) $\alpha$-leakage in privacy-preserving data publishing [15].
While the Sibson and Arimoto MI have closed-form expressions, the Augustin–Csiszár MI, Lapidoth–Pfister MI, and $\alpha$-capacity generally do not; therefore, these types of MI and $\alpha$-capacity must be calculated by solving optimization problems. Algorithms for computing the $\alpha$-MI and $\alpha$-capacity have been explored [2, 16, 17, 3, 18, 19, 20]. For the Augustin–Csiszár MI for $\alpha\in(0,1)$, Augustin [3] proposed an iterative algorithm, which Karakos et al. [18] rediscovered as an alternating optimization (AO) algorithm. Recently, Wang et al. [19] proposed an algorithm based on geodesically convex optimization, which is effective for $\alpha\in(1,\infty)$. For the Sibson and Arimoto capacities, Arimoto [2, 16] proposed AO algorithms similar to the Arimoto–Blahut algorithm [21, 22] for calculating the channel capacity $C$. Recently, Kamatsuka et al. [17] proposed additional AO algorithms for computing these capacities and demonstrated that they are equivalent to those suggested by Arimoto. For Augustin–Csiszár capacity, Kamatsuka et al. [20] proposed an AO algorithm. Table I summarizes the $\alpha$-MI, $\alpha$-capacity, and their calculation methods.
TABLE I: Examples of $\alpha$-MI and their calculation methods
Name Calculation method
Sibson MI
$I_{\alpha}^{\text{S}}(X;Y)$
Sibson capacity
$C_{\alpha}^{\text{S}}$
Closed-form expression
AO algorithms [2, 17], (This work)
Arimoto MI
$I_{\alpha}^{\text{A}}(X;Y)$
Arimoto capacity
$C_{\alpha}^{\text{A}}$
Closed-form expression
AO algorithms [16, 17]
Augustin–Csiszár MI
$I_{\alpha}^{\text{C}}(X;Y)$
Augustin–Csiszár capacity
$C_{\alpha}^{\text{C}}$
AO algorithm for $\alpha\in(0,1)$ [3, 18]
Geodesically convex optimization [19]
AO algorithm for $\alpha\in(1,\infty)$ (This work)
AO algorithm for $\alpha\in(1,\infty)$ [20]
Lapidoth–Pfister MI
$I_{\alpha}^{\text{LP}}(X;Y)$ [5], [6]
Lapidoth–Pfister capacity
$C_{\alpha}^{\text{LP}}$
AO algorithms (This work)
AO algorithm for $\alpha\in(1,\infty)$ (This work)
These AO algorithms are based on the variational characterization of $\alpha$-MI, transforming the definitions of $\alpha$-MI into optimization problems. This study proposes novel variational characterizations of $\alpha$-MI and presents AO algorithms based on these characterizations. The main contributions of this paper are as follows:
  • Extending the results of the variational characterization of the Sibson MI by Shayevitz [23], we provide a novel proof for Arimoto’s variational characterization of the Sibson MI using the reverse channel $r_{X\mid Y}$ [2] (Theorem 1). As a byproduct of this proof, we also present a new characterization of the Sibson MI. Similarly, we provide novel variational characterizations of the Augustin–Csiszár and Lapidoth–Pfister MI (Theorems 2 and 3).
  • Based on the characterizations, we propose novel AO algorithms for computing Augustin–Csiszár MI, Lapidoth–Pfister MI, and Sibson and Lapidoth–Pfister capacities (Section IV).
  • We compare the performance of the AO algorithms for computing $\alpha$-capacity through numerical examples (Section V).
II Preliminaries
Let $(X,Y)$ denote random variables distributed according to a joint distribution $p_{X,Y}=p_{X}p_{Y\mid X}$ on finite alphabets $\mathcal{X}\times\mathcal{Y}$. Let $H(X):=-\sum_{x}p_{X}(x)\log p_{X}(x)$ be the Shannon entropy, $H(X|Y):=-\sum_{y}p_{Y}(y)\sum_{x}p_{X\mid Y}(x|y)\log p_{X\mid Y}(x|y)$ the conditional entropy, and $I(X;Y):=H(X)-H(X|Y)$ the Shannon MI, where $p_{Y}$ is the marginal distribution of $Y$. We denote the expectation of $f(X)$ as $\mathbb{E}_{X}^{p_{X}}\left[f(X)\right]:=\sum_{x}p_{X}(x)f(x)$. Throughout this paper, we use $\log$ to represent the natural logarithm.
We initially review the Shannon MI and $\alpha$-MI. The Shannon MI $I(X;Y)=I(p_{X},p_{Y\mid X})$ have the following variational characterizations.
Proposition 1 ([24, Cor 4.2, Thm 4.3 and 4.4]).
$I(p_{X},p_{Y\mid X})$ $=\min_{q_{Y}}D(p_{X}p_{Y\mid X}||p_{X}q_{Y})$ (1)
$=\min_{q_{X}}\min_{q_{Y}}D(p_{X}p_{Y\mid X}||q_{X}q_{Y})$ (2)
$=\max_{r_{X\mid Y}}\mathbb{E}_{X,Y}^{p_{X}p_{Y\mid X}}\log\left[\frac{r_{X\mid Y}(X\mid Y)}{p_{X}(X)}\right],$ (3)
where the minimum in (1) is taken over all distributions on $\mathcal{Y}$ and is achieved at $q_{Y}^{*}=p_{Y}$. The minimum in (2) is taken over all product distributions on $\mathcal{X}\times\mathcal{Y}$ and is achieved at $(q_{X}^{*},q_{Y}^{*})=(p_{X},p_{Y})$. The maximum in (3) is taken over all the family of conditional distributions, i.e., reverse channels $r_{X\mid Y}:=\{r_{X\mid Y}(\cdot|y)\}_{y\in\mathcal{Y}}$ and is achieved at $r^{*}_{X\mid Y}=p_{X\mid Y}$.
Definition 1.
Let $\alpha\in(0,1)\cup(1,\infty)$ and $(X,Y)\sim p_{X,Y}=p_{X}p_{Y\mid X}$. Then, the Sibson MI of order $\alpha$, Arimoto MI of order $\alpha$, Augustin–Csiszár MI of order $\alpha$, and Lapidoth–Pfister MI of order $\alpha$ are defined as follows, respectively:
$I_{\alpha}^{\text{S}}(X;Y)$ $:=\min_{q_{Y}}D_{\alpha}(p_{X}p_{Y\mid X}||p_{X}q_{Y}),$ (4)
$I_{\alpha}^{\text{A}}(X;Y)$ $:=H_{\alpha}(X)-H_{\alpha}^{\text{A}}(X\mid Y),$ (5)
$I_{\alpha}^{\text{C}}(X;Y)$ $:=\min_{q_{Y}}\mathbb{E}_{X}^{p_{X}}\left[D_{\alpha}(p_{Y\mid X}(\cdot\mid X)||q_{Y})\right],$ (6)
$I_{\alpha}^{\text{LP}}(X;Y)$ $:=\min_{q_{X}}\min_{q_{Y}}D_{\alpha}(p_{X}p_{Y\mid X}||q_{X}q_{Y}),$ (7)
where
$D_{\alpha}(p||q):=\frac{1}{\alpha-1}\log\sum_{z}p(z)^{\alpha}q(z)^{1-\alpha}$ (8)
is the Rényi divergence between $p$ and $q$ of order $\alpha$, $H_{\alpha}(X):=\frac{1}{1-\alpha}\log\sum_{x}p_{X}(x)^{\alpha}$ is the Rényi entropy of order $\alpha$, and $H_{\alpha}^{\text{A}}(X|Y):=\frac{\alpha}{1-\alpha}\log\sum_{y}\left(\sum_{x}p_{X}(x)^{\alpha}p_{Y\mid X}(y|x)^{\alpha}\right)^{\frac{1}{\alpha}}$ is the Arimoto conditional entropy of order $\alpha$ [2].
Remark 1.
The Sibson and Arimoto MI have the following closed-form expressions using the Gallager error exponent function $E_{0}(\rho,p_{X}):=-\log\sum_{y}\left(\sum_{x}p_{X}(x)p_{Y\mid X}(y|x)^{\frac{1}{1+\rho}}\right)^{1+\rho}$ [9]:
$I_{\alpha}^{\text{S}}(X;Y)$ $=\frac{\alpha}{1-\alpha}E_{0}\left(\frac{1}{\alpha}-1,p_{X}\right),$ (9)
$I_{\alpha}^{\text{A}}(X;Y)$ $=\frac{\alpha}{1-\alpha}E_{0}\left(\frac{1}{\alpha}-1,p_{X_{\alpha}}\right),$ (10)
where $p_{X_{\alpha}}(x):=\frac{p_{X}(x)^{\alpha}}{\sum_{x}p_{X}(x)^{\alpha}}$ is the $\alpha$-tilted distribution [15] (also known as the escort distribution [25]) of $p_{X}$. However, the closed-form expressions of the Augustin–Csiszár MI and Lapidoth–Pfister MI are unknown and must be computed by solving the optimization problems.
Karakos et al. [18], Shayevitz and Ofer [23], and Lapidoth and Pfister [5] provided the following variational characterizations of $\alpha$-MI.
Proposition 2 ([18], [23, Thm 1], [5, Lemma 7]).
Let $\alpha\in(0,1)\cup(1,\infty)$. Then,
$(1-\alpha)I_{\alpha}^{\text{C}}(X;Y)$
$=\min_{\tilde{q}_{Y\mid X}}\left\{\alpha D(p_{X}\tilde{q}_{Y\mid X}||p_{X}p_{Y\mid X})+(1-\alpha)I(p_{X},\tilde{q}_{Y\mid X})\right\},$ (11)
$(1-\alpha)I_{\alpha}^{\text{S}}(X;Y)$
$=\min_{\tilde{q}_{X,Y}}\{\alpha D(p_{X}\tilde{q}_{Y\mid X}||p_{X}p_{Y\mid X})$
$\qquad\qquad\qquad+(1-\alpha)I(p_{X},\tilde{q}_{Y\mid X})+D(\tilde{q}_{X}||p_{X})\},$ (12)
$(1-\alpha)I_{\alpha}^{\text{LP}}(X;Y)=\min_{\tilde{q}_{X,Y}}\{\alpha D(\tilde{q}_{X,Y}||p_{X}p_{Y\mid X})$
$\qquad\qquad\qquad\qquad\qquad+(1-\alpha)I(\tilde{q}_{X},\tilde{q}_{Y\mid X})\}.$ (13)
III Variational Characterizations of $\alpha$-MI
It is natural to consider variational characterizations of $\alpha$-MI similar to the Shannon MI characterization in Proposition 1, some of which were already obtained by Arimoto [2, 16], Kamatsuka et al. [17], and Karakos et al. [18]. In this section, we derive new characterizations of Sibson, Augustin–Csiszár, and Lapidoth–Pfister MI through a novel proof of Arimoto’s characterization.
Theorem 1 ([2, Thm 4],[17, Thm 1]).
Let $\alpha\in(0,1)\cup(1,\infty)$. Then,
$I_{\alpha}^{\text{S}}(X;Y)$ $=\max_{r_{X\mid Y}}F_{\alpha}^{\text{S1}}(p_{X},r_{X\mid Y})$ (14)
$=\max_{r_{X\mid Y}}F_{\alpha}^{\text{S2}}(p_{X},r_{X\mid Y}),$ (15)
where
$F_{\alpha}^{\text{S1}}(p_{X},r_{X\mid Y})$
$:=\frac{\alpha}{\alpha-1}\log\sum_{x,y}p_{X}(x)^{\frac{1}{\alpha}}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}},$ (16)
$F_{\alpha}^{\text{S2}}(p_{X},r_{X\mid Y}):=F_{\alpha}^{\text{S1}}(p_{X},r_{X_{\alpha}\mid Y})$
$=\frac{\alpha}{\alpha-1}\log\sum_{x,y}p_{X}(x)^{\frac{1}{\alpha}}p_{Y\mid X}(y|x)r_{X_{\alpha}\mid Y}(x|y)^{1-\frac{1}{\alpha}}.$ (17)
For $\alpha\in(1,\infty)$,
$I_{\alpha}^{\text{S}}(X;Y)$ $=\max_{\tilde{q}_{X,Y}}\max_{r_{X\mid Y}}\tilde{F}_{\alpha}^{\text{S3}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y}),$ (18)
where
$\tilde{F}_{\alpha}^{\text{S3}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y}):=\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||\tilde{q}_{X}p_{Y\mid X})$
$+\mathbb{E}_{X,Y}^{\tilde{q}_{X,Y}}\left[\log\frac{r_{X\mid Y}(X|Y)}{\tilde{q}_{X}(X)}\right]+\frac{1}{1-\alpha}D(\tilde{q}_{X}||p_{X}).$ (19)
Proof.
See Appendix A. ∎
Remark 2.
From (9) and (10), similar characterizations of the Arimoto MI can be obtained by replacing $p_{X}$ in (14), (15), and (18) with $p_{X_{\alpha}}$, two of which were first obtained by Arimoto [16] and Kamatsuka et al. [17].
Similar to the proof for the Sibson MI, the following variational characterizations of the Augustin–Csiszár MI and Lapidoth–Pfister MI are obtained.
Theorem 2.
Let $\alpha\in(0,1)\cup(1,\infty)$. Then,
$I_{\alpha}^{\text{C}}(X;Y)$
=$=\begin{cases}\min_{\tilde{q}_{Y\mid X}}\min_{q_{Y}}F_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X},q_{Y}),&\alpha\in(0,1),\\ \max_{\tilde{q}_{Y\mid X}}\max_{r_{X\mid Y}}\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X},r_{X\mid Y}),&\alpha\in(1,\infty)\end{cases}$ (20)
$=H(p_{X})$
$+\max_{r_{X\mid Y}}\frac{\alpha}{\alpha-1}\sum_{x}p_{X}(x)\log\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}},$ (21)
where
${F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X},q_{Y})$
$=\frac{\alpha}{1-\alpha}D(p_{X}\tilde{q}_{Y\mid X}||p_{X}p_{Y\mid X})+D(p_{X}\tilde{q}_{Y\mid X}||p_{X}q_{Y})$ (22)
$\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X},r_{X\mid Y})$
$:=\frac{\alpha}{1-\alpha}D(p_{X}\tilde{q}_{Y\mid X}||p_{X}p_{Y\mid X})+\mathbb{E}_{X,Y}^{p_{X}\tilde{q}_{Y\mid X}}\left[\frac{r_{X\mid Y}(X\mid Y)}{p_{X}(X)}\right].$ (23)
Proof.
See Appendix B. ∎
Theorem 3.
Let $\alpha\in(0,1)\cup(1,\infty)$. Then,
$I_{\alpha}^{\text{LP}}(X;Y)$
=$=\begin{cases}\min_{\tilde{q}_{X,Y}}\min_{q_{Y}}F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},q_{Y}),&\alpha\in(0,1),\\ \max_{\tilde{q}_{X,Y}}\max_{r_{X\mid Y}}\tilde{F}_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},r_{X\mid Y}),&\alpha\in(1,\infty),\end{cases}$ (24)
where $\tilde{q}_{X,Y}=\tilde{q}_{X}\tilde{q}_{Y\mid X}$ and
$F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},q_{Y}):=\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||\tilde{q}_{X}p_{Y\mid X})$
$+D(\tilde{q}_{X}\tilde{q}_{Y\mid X}||\tilde{q}_{X}q_{Y})+\frac{\alpha}{1-\alpha}D(\tilde{q}_{X}||p_{X}),$ (25)
$\tilde{F}_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},r_{X\mid Y}):=\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||\tilde{q}_{X}p_{Y\mid X})$
$+\mathbb{E}_{X,Y}^{\tilde{q}_{X,Y}}\left[\log\frac{r_{X\mid Y}(X|Y)}{\tilde{q}_{X}(X)}\right]+\frac{\alpha}{1-\alpha}D(\tilde{q}_{X}||p_{X}).$ (26)
For $\alpha\in(1,\infty)$,
$I_{\alpha}^{\text{LP}}(X;Y)=\max_{r_{X\mid Y}}\frac{2\alpha-1}{\alpha-1}$
×$\times\log\sum_{x}p_{X}(x)^{\frac{\alpha}{2\alpha-1}}\left(\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}\right)^{\frac{\alpha}{2\alpha-1}}.$ (27)
Proof.
See Appendix C. ∎
IV AO Algorithms
This section proposes novel AO algorithms for computing Augustin–Csiszár MI, Lapidoth–Pfister MI, and $\alpha$-capacities based on the variational characterizations obtained in Section III.
IV-A AO Algorithms for Computing the Augustin–Csiszár MI
Karakos et al. [18] proposed an AO algorithm based on the variational characterization in (20) for the Augustin–Csiszár MI when $\alpha\in(0,1)$. We propose an AO algorithm for $\alpha\in(1,\infty)$ with the following updating formulae:
Proposition 3.
Let $\alpha\in(1,\infty)$. Then, the following holds:
  1. 1.
    For a fixed $\tilde{q}_{Y\mid X}$, $\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X},r_{X\mid Y})$ is maximized by
    $r_{X\mid Y}^{*}(x\mid y)$ $:=\frac{p_{X}(x)\tilde{q}_{Y\mid X}(y\mid x)}{\sum_{x}p_{X}(x)\tilde{q}_{Y\mid X}(y\mid x)}.$ (28)
  2. 2.
    For a fixed $r_{X\mid Y}$, $\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X},r_{X\mid Y})$ is maximized by
    $\tilde{q}_{Y\mid X}^{*}(y\mid x)$ $:=\frac{p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}}{\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}}.$ (29)
Proof.
This follows from the proof of Theorem 2 and Proposition 1. ∎
IV-B AO Algorithms for Computing the Lapidoth–Pfister MI
The Lapidoth–Pfister MI in (7) is defined as a double minimization problem; $I_{\alpha}^{\text{LP}}(X;Y):=\min_{q_{X}}\min_{q_{Y}}D_{\alpha}(p_{X}p_{Y\mid X}||q_{X}q_{Y})$. From the definition and properties of the Lapidoth–Pfister MI ([5, Prop 9 and Lemma 16], [6, Lemma 29]), we derive an AO algorithm for the Lapidoth–Pfister MI with the following updating formulae.
Corollary 1.
Let $\alpha\in(0,1)\cup(1,\infty)$ and $p_{X,Y}=p_{X}p_{Y\mid X}$. Then, the following holds:
  1. 1.
    For a fixed $q_{Y}$, $D_{\alpha}(p_{X}p_{Y\mid X}||q_{X}q_{Y})$ is minimized by
    $q_{X}^{*}(x)$ $:=\frac{\left[\sum_{y}p_{X,Y}(x,y)^{\alpha}q_{Y}(y)^{1-\alpha}\right]^{\frac{1}{\alpha}}}{\sum_{x}\left[\sum_{y}p_{X,Y}(x,y)^{\alpha}q_{Y}(y)^{1-\alpha}\right]^{\frac{1}{\alpha}}}.$ (30)
  2. 2.
    For a fixed $q_{X}$, $D_{\alpha}(p_{X}p_{Y\mid X}||q_{X}q_{Y})$ is minimized by
    $q_{Y}^{*}(y)$ $:=\frac{\left[\sum_{x}p_{X,Y}(x,y)^{\alpha}q_{X}(x)^{1-\alpha}\right]^{\frac{1}{\alpha}}}{\sum_{y}\left[\sum_{x}p_{X,Y}(x,y)^{\alpha}q_{X}(x)^{1-\alpha}\right]^{\frac{1}{\alpha}}}.$ (31)
Additional AO algorithm can be derived based on the Lapidoth–Pfister MI characterization in (25).
Proposition 4.
Let $F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},q_{Y})$ and $\tilde{F}_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},r_{X\mid Y})$ be functionals defined in (25) and (26), respectively.
Then, for $\alpha\in(0,1)$, the following holds:
  1. 1.
    For a fixed $\tilde{q}_{X,Y}$, $F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},q_{Y})$ is minimized by
    $q_{Y}^{*}(y)$ $:=\sum_{x}\tilde{q}_{X,Y}(x,y).$ (32)
  2. 2.
    For a fixed $q_{Y}$, $F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},q_{Y})$ is minimized by
    $\tilde{q}_{X,Y}^{*}(x,y)$ $=\bar{q}_{X}(x)\bar{q}_{Y\mid X}(y\mid x),$ (33)
    where
    $\bar{q}_{X}(x):=\frac{p_{X}(x)\left(\sum_{y}p_{Y\mid X}(y\mid x)^{\alpha}q_{Y}(y)^{1-\alpha}\right)^{\frac{1}{\alpha}}}{\sum_{x}p_{X}(x)\left(\sum_{y}p_{Y\mid X}(y\mid x)^{\alpha}q_{Y}(y)^{1-\alpha}\right)^{\frac{1}{\alpha}}},$ (34)
    $\bar{q}_{Y\mid X}(y\mid x):=\frac{p_{Y\mid X}(y\mid x)^{\alpha}q_{Y}(y)^{1-\alpha}}{\sum_{y}p_{Y\mid X}(y\mid x)^{\alpha}q_{Y}(y)^{1-\alpha}}.$ (35)
For $\alpha\in(1,\infty)$, the following holds:
  1. 3)
    For a fixed $\tilde{q}_{X,Y}$, $\tilde{F}_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},r_{X\mid Y})$ is maximized by
    $r_{X\mid Y}^{*}(x\mid y)$ $=\frac{\tilde{q}_{X,Y}(x,y)}{\sum_{x}\tilde{q}_{X,Y}(x,y)}.$ (36)
  2. 4)
    For a fixed $r_{X\mid Y}$, $\tilde{F}_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},r_{X\mid Y})$ is maximized by
    $q_{X,Y}^{*}(x,y)$ $:=\hat{q}_{X}(x)\hat{q}_{Y\mid X}(y\mid x),$ (37)
    where
    $\hat{q}_{X}(x)$
    $:=\frac{p_{X}(x)^{\frac{\alpha}{2\alpha-1}}\left(\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}\right)^{\frac{\alpha}{2\alpha-1}}}{\sum_{x}p_{X}(x)^{\frac{\alpha}{2\alpha-1}}\left(\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}\right)^{\frac{\alpha}{2\alpha-1}}},$ (38)
    $\hat{q}_{Y\mid X}(y\mid x):=\frac{p_{Y\mid X}(y\mid x)r_{X\mid Y}(x\mid y)^{1-\frac{1}{\alpha}}}{\sum_{x}p_{Y\mid X}(y\mid x)r_{X\mid Y}(x\mid y)^{1-\frac{1}{\alpha}}}.$ (39)
Proof.
See Appendix D. ∎
IV-C AO Algorithms for Computing $\alpha$-Capacity
Arimoto [2] proposed an AO algorithm for computing the Sibson capacity based on the characterization in (14). Subsequently, Arimoto [16] and Kamatsuka et al. [17] proposed similar AO algorithms, building on the characterizations of Arimoto and Sibson MI in Remark 2 and (15) (see [17, Thm 1 and Table I]). Kamatsuka et al. [20] derived an AO algorithm for computing the Augustin–Csiszár capacity for $\alpha\in(1,\infty)$ based on the characterization in (20). In [20], they also proposed another AO algorithm for the Sibson capacity by modifying the AO algorithm for computing the correct decoding exponent proposed by Jitsumatsu and Oohama [26, Algorithm 1].
In addition to these algorithms, an AO algorithm for $\alpha\in(1,\infty)$ can be derived from the characterizations in (18) and its proof with the following updating formulae:
Proposition 5.
Let $\alpha\in(1,\infty)$ and $\tilde{F}_{\alpha}^{\text{S3}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y})$ be a functional defined in (19). Then,
  1. 1.
    For a fixed $(p_{X},\tilde{q}_{X,Y})$, $\tilde{F}_{\alpha}^{\text{S3}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y})$ is maximized by
    $r_{X\mid Y}^{*}(x\mid y)$ $:=\frac{\tilde{q}_{X,Y}(x,y)}{\sum_{x}\tilde{q}_{X,Y}(x,y)}.$ (40)
  2. 2.
    For a fixed $(\tilde{q}_{X,Y},r_{X\mid Y})$, $\tilde{F}_{\alpha}^{\text{S3}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y})$ is maximized by
    $p_{X}^{*}(x):=\sum_{y}\tilde{q}_{X,Y}(x,y).$ (41)
  3. 3.
    For a fixed $(p_{X},r_{X\mid Y})$, $\tilde{F}_{\alpha}^{\text{S3}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y})$ is maximized by
    $\tilde{q}^{*}_{X,Y}(x,y)$ $:=\frac{p_{X}(x)^{\frac{1}{\alpha}}p_{Y\mid X}(y\mid x)r_{X\mid Y}(x\mid y)^{1-\frac{1}{\alpha}}}{\sum_{x,y}p_{X}(x)^{\frac{1}{\alpha}}p_{Y\mid X}(y\mid x)r_{X\mid Y}(x\mid y)^{1-\frac{1}{\alpha}}}.$ (42)
Remark 3.
The AO algorithm with the above updating formulae is equivalent to the AO algorithm for computing the correct decoding exponent without a cost constraint in [26, Algorithm 2].
Similarly, a novel AO algorithm for computing the Lapidoth–Pfister capacity for $\alpha\in(1,\infty)$ can be derived from the characterization of Lapidoth–Pfister MI in (24) and its proof using the following updating formulae:
Proposition 6.
Let $\alpha\in(1,\infty)$ and $\tilde{F}_{\alpha}^{\text{LP}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y})$ be a functional defined in (26) [^1] Here, we have added $p_{X}$ as an argument for the functional $\tilde{F}_{\alpha}^{\text{LP}}$.. Then,
  1. 1.
    For a fixed $(p_{X},\tilde{q}_{X,Y})$, $\tilde{F}_{\alpha}^{\text{LP}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y})$ is maximized by
    $r_{X\mid Y}^{*}(x\mid y)$ $:=\frac{\tilde{q}_{X,Y}(x,y)}{\sum_{x}\tilde{q}_{X,Y}(x,y)}.$ (43)
  2. 2.
    For a fixed $(\tilde{q}_{X,Y},r_{X\mid Y})$, $\tilde{F}_{\alpha}^{\text{LP}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y})$ is maximized by
    $p_{X}^{*}(x):=\sum_{y}\tilde{q}_{X,Y}(x,y).$ (44)
  3. 3.
    For a fixed $(p_{X},r_{X\mid Y})$, $\tilde{F}_{\alpha}^{\text{LP}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y})$ is maximized by
    $\tilde{q}^{*}_{X,Y}(x,y)$ $:=\hat{q}_{X}(x)\hat{q}_{Y\mid X}(y\mid x),$ (45)
    where $\hat{q}_{X}$ and $\hat{q}_{Y\mid X}$ are defined in (38) and (39), respectively.
Proof.
This follows from Proposition 1 and the proof of Proposition 4. ∎
IV-D Global Convergence Properties of AO Algorithms
This subsection discusses the global convergence properties of each AO algorithm.
IV-D1 AO algorithms for computing $\alpha$-capacity
Arimoto [27, Thm 3] and Kamatsuka et al. [17, Cor 2] demonstrated the global convergence properties of the AO algorithms based on the characterization of the Sibson MI in (14) and (15) as well as the corresponding AO algorithms derived from the characterization of the Arimoto MI. The modified Jitsuamtsu–Oohama algorithm proposed in [20] also has this property (see [26, Thm 2]), while the property of the AO algorithm based on the characterization in (20) for computing the Augustin–Csiszár capacity for $\alpha\in(1,\infty)$ remains undemonstrated.
Here, we prove that the AO algorithm for computing the Lapidoth–Pfister capacity for $\alpha\in(1,\infty)$ in Proposition 6 has the global convergence property. First, the following lemma follows directly from Theorem 3 and Proposition 6.
Lemma 1.
Let $\alpha\in(1,\infty)$ and $\{p_{X}^{(k)}\}_{k=0}^{\infty}$, $\{q_{X,Y}^{(k)}\}_{k=0}^{\infty}$, and $\{r_{X\mid Y}^{(k)}\}_{k=0}^{\infty}$ be sequences of distributions at the $k$th iteration of the AO algorithm. Then, the following holds:
$\tilde{F}_{\alpha}^{\text{LP}}(p_{X}^{(0)},\tilde{q}_{X,Y}^{(0)},r_{X\mid Y}^{(0)})\leq\cdots\leq\tilde{F}_{\alpha}^{\text{LP}}(p_{X}^{(k)},\tilde{q}_{X,Y}^{(k)},r_{X\mid Y}^{(k)})$
$\leq\tilde{F}_{\alpha}^{\text{LP}}(p_{X}^{(k+1)},\tilde{q}_{X,Y}^{(k+1)},r_{X\mid Y}^{(k)})\leq\cdots\leq C_{\alpha}^{\text{LP}}.$ (46)
Next, the approximation error at the $k$th iteration, $C_{\alpha}^{\text{LP}}-\tilde{F}_{\alpha}^{\text{C}}(p_{X}^{(k+1)},\tilde{q}_{X,Y}^{(k+1)},r_{X\mid Y}^{(k)})$, can be upper bounded as follows:
Lemma 2.
Let $\alpha\in(1,\infty)$. Then, the approximation error has the following upper bound:
$C_{\alpha}^{\text{LP}}-\tilde{F}_{\alpha}^{\text{LP}}(p_{X}^{(k+1)},\tilde{q}_{X,Y}^{(k+1)},r_{X\mid Y}^{(k)})\leq\frac{\alpha}{\alpha-1}\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)$
×$\times\log\frac{\tilde{q}_{X,Y}^{(k+1)}(x,y)}{\tilde{q}_{X,Y}^{(k)}(x,y)}+\sum_{x}\tilde{q}_{X}^{\text{opt}}(x)\log\frac{\tilde{q}_{X}^{(k+1)}(x)}{\tilde{q}_{X}^{(k)}(x)},$ (47)
where $\tilde{q}_{X,Y}^{\text{opt}}$ is the optimal joint distribution that achieves the maximum in (24).
Proof.
See Appendix E. ∎
Summing (47) from $k=0$ to $k=N$ obtains the following upper bound that is independent of $N$:
$\sum_{k=0}^{N}\left(C_{\alpha}^{\text{LP}}-\tilde{F}_{\alpha}^{\text{LP}}(p_{X}^{(k+1)},\tilde{q}_{X,Y}^{(k+1)}r_{X\mid Y}^{(k)})\right)$ (48)
$\leq\frac{\alpha}{\alpha-1}\sum_{x,y}\tilde{q}_{X,Y}^{*}(x,y)\log\frac{\tilde{q}_{X,Y}^{(N+1)}(x,y)}{\tilde{q}_{X,Y}^{(0)}(x,y)}$
$\qquad\qquad\qquad+\sum_{x}\tilde{q}_{X}^{*}(x)\log\frac{\tilde{q}_{X}^{(N+1)}(x)}{\tilde{q}_{X}^{(0)}(x)}$ (49)
$=\frac{\alpha}{\alpha-1}\left(D(\tilde{q}_{X,Y}^{*}||\tilde{q}_{X,Y}^{(0)})-D(\tilde{q}_{X,Y}^{*}||\tilde{q}_{X,Y}^{(N+1)})\right)$
$\qquad\qquad\qquad+D(\tilde{q}_{X}^{*}||\tilde{q}_{X}^{(0)})-D(\tilde{q}_{X}^{*}||\tilde{q}_{X}^{(N+1)})$ (50)
$\leq\frac{\alpha}{\alpha-1}D(\tilde{q}_{X,Y}^{*}||\tilde{q}_{X,Y}^{(0)})+D(\tilde{q}_{X}^{*}||\tilde{q}_{X}^{(0)}).$ (51)
If we combine this upper bound with Lemma 1, then the infinite series $\sum_{k=0}^{\infty}\left(C_{\alpha}^{\text{LP}}-\tilde{F}_{\alpha}^{\text{LP}}(p_{X}^{(k+1)},\tilde{q}_{X,Y}^{(k+1)},r_{X\mid Y}^{(k)})\right)$ converges. Therefore, $C_{\alpha}^{\text{LP}}-\tilde{F}_{\alpha}^{\text{LP}}(p_{X}^{(k+1)},\tilde{q}_{X,Y}^{(k+1)},r_{X\mid Y}^{(k)})\to 0$ as $k\to\infty$, which implies the convergence result of the Lapidoth–Pfister capacity.
Corollary 2.
Let $\alpha\in(1,\infty)$. Then,
$\lim_{k\to\infty}\tilde{F}_{\alpha}^{\text{LP}}(p_{X}^{(k+1)},\tilde{q}_{X,Y}^{(k+1)},r_{X\mid Y}^{(k)})=C_{\alpha}^{\text{LP}}.$ (52)
IV-D2 AO algorithm for computing Augustin–Csiszár MI
Karakos et al. [18, Thm 1] proved that the AO algorithm based on (20) and (22) converges to the Augustin–Csiszár MI for $\alpha\in(0,1)$. Similar to the proof of convergence of the Lapidoth–Pfister capacity, we show that the AO algorithm based on (20) and (23) globally converges when $\alpha\in(1,\infty)$. Hence, it suffices to prove the following lemmas:
Lemma 3.
Let $\alpha\in(1,\infty)$ and $\{\tilde{q}_{Y\mid X}^{(k)}\}_{k=0}^{\infty}$ and $\{r_{X\mid Y}^{(k)}\}_{k=0}^{\infty}$ be sequences of distributions at the $k$th iteration of the AO algorithm. Then, the following holds:
$\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X}^{(0)},r_{X\mid Y}^{(0)})\leq\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X}^{(1)},r_{X\mid Y}^{(0)})\leq\cdots$
$\leq\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X}^{(k)},r_{X\mid Y}^{(k)})\leq\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X}^{(k+1)},r_{X\mid Y}^{(k)})\leq\cdots\leq I_{\alpha}^{\text{C}}(X;Y).$ (53)
Proof.
This follows directly from Theorem 2 and Proposition 3. ∎
Lemma 4.
Let $\alpha\in(1,\infty)$. Then, the approximation error has the following upper bound:
$I_{\alpha}^{\text{C}}(X;Y)-\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X}^{(k+1)},r_{X\mid Y}^{(k)})$
$\leq\frac{\alpha}{\alpha-1}\sum_{x,y}p_{X}(x)\tilde{q}_{Y\mid X}^{\text{opt}}(y\mid x)\log\frac{\tilde{q}_{Y\mid X}^{(k+1)}(y\mid x)}{\tilde{q}_{Y\mid X}^{(k)}(y\mid x)},$ (54)
where $\tilde{q}_{Y\mid X}^{\text{opt}}=\{\tilde{q}_{Y\mid X}^{\text{opt}}(\cdot|x)\}_{x\in\mathcal{X}}$ is the optimal family of distributions that achieves the maximum in (20)
Proof.
See Appendix F. ∎
The following convergence result can then be obtained:
Corollary 3.
Let $\alpha\in(1,\infty)$. Then,
$\lim_{k\to\infty}\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X}^{(k+1)},r_{X\mid Y}^{(k)})$ $=I_{\alpha}^{\text{C}}(X;Y).$ (55)
IV-D3 AO algorithms for computing Lapidoth–Pfister MI
The global convergence property of the AO algorithm based on Corollary 1 follows immediately from the sufficient condition for the global convergence of general AO algorithms [28, Thm 10.5].
Proposition 7.
Let $\alpha\in[1/2,1)\cup(1,\infty)$ and $\{q_{X}^{(k)}\}_{k=0}^{\infty}$ and $\{q_{Y}^{(k)})\}_{k=0}^{\infty}$ be sequences of distributions at the $k$th iteration of the AO algorithm. Then,
$\lim_{k\to\infty}D_{\alpha}(p_{X}p_{Y\mid X}||q_{X}^{(k)}q_{Y}^{(k)})$ $=I_{\alpha}^{\text{LP}}(X;Y).$ (56)
Proof.
We can prove (56) by combining the sufficient condition from [28, Thm 10.5] with the fact that $(q_{X},q_{Y})\mapsto D_{\alpha}(p_{X}p_{Y\mid X}||q_{X}q_{Y})$ is jointly convex for $[1/2,1)\cup(1,\infty)$ [5, Lemma 15]. ∎
Remark 4.
Recently, Tsai et al. [29] provided a non-asymptotic analysis of the linear convergence rate of the AO algorithm.
The global convergence property of the AO algorithm based on (24) for $\alpha\in(1,\infty)$ directly follows from Corollary 2 by replacing $p_{X}^{(k)}$ with $p_{X}$ for all $k\in\mathbb{N}$. For $\alpha\in(0,1)$, it can be shown similarly to Corollaries 2 and 3; thus, it suffices to show the following lemmas:
Lemma 5.
Let $\alpha\in(0,1)$ and $\{\tilde{q}_{X,Y}^{(k)}\}_{k=0}^{\infty}$ and $\{q_{Y}^{(k)}\}_{k=0}^{\infty}$ be sequences of distributions at the $k$th iteration of the AO algorithm. Then, the following holds:
$F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y}^{(0)},q_{Y}^{(0)})\geq F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y}^{(1)},q_{Y}^{(0)})\geq\cdots$
$\geq F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y}^{(k)},q_{Y}^{(k)})\geq F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y}^{(k+1)},q_{Y}^{(k)})\geq\cdots\geq I_{\alpha}^{\text{LP}}(X;Y).$ (57)
Proof.
This follows directly from Theorem 2 and Proposition 3. ∎
Lemma 6.
Let $\alpha\in[1/2,1)$. Then, the approximation error has the following upper bound:
$F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y}^{(k+1)},q_{Y}^{(k)})-I_{\alpha}^{\text{LP}}(X;Y)$
$\qquad\leq\frac{\alpha}{1-\alpha}\sum_{x}\tilde{q}_{X}^{\text{opt}}(x)\log\frac{\tilde{q}_{X}^{(k+1)}(x)}{\tilde{q}_{X}^{(k)}(x)},$ (58)
where $\tilde{q}_{X}^{\text{opt}}(x):=\sum_{y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)$ and $\tilde{q}_{X,Y}^{\text{opt}}$ is the optimal joint distributions that achieve the minimum in (24).
Proof.
See Appendix G. ∎
Corollary 4.
Let $\alpha\in[1/2,1)\cup(1,\infty)$. Then,
$\lim_{k\to\infty}F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y}^{(k+1)},q_{Y}^{(k)})$ $=I_{\alpha}^{\text{LP}}(X;Y),\quad\alpha\in[1/2,1),$ (59)
$\lim_{k\to\infty}\tilde{F}_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y}^{(k+1)},r_{X\mid Y}^{(k)})$ $=I_{\alpha}^{\text{LP}}(X;Y),\quad\alpha\in(1,\infty).$ (60)
V Comparison of AO algorithms for $\alpha$-Capacity
It is known that the Sibson, Arimoto, and Augustin–Csiszár capacity are equivalent: $C_{\alpha}^{\text{S}}=C_{\alpha}^{\text{A}}=C_{\alpha}^{\text{C}}$ for $\alpha\in(0,1)\cup(1,\infty)$ [2],[4],[30]. Moreover, the Lapidoth–Pfister capacity $C_{\alpha}^{\text{LP}}$ is equivalent to these capacities for $\alpha\in(1,\infty)$ [30, Thm 4], [31, p.4]. Thus, all AO algorithms obtained so far compute equivalent quantities. This section compares their performance through numerical examples. From this point onward, we refer to an AO algorithm based on the variational characterization using a functional $F_{\alpha}^{\text{X}}(\cdot)$ or $\tilde{F}_{\alpha}^{\text{X}}(\cdot)$ as Algorithm X. Moreover, we refer to the modified AO algorithm of [26, Algorithm 1] as Algorithm JO.
Kamatsuka et al. [17, Cor 1] proved that the AO algorithms S1, S2, and the algorithms based on the variational characterizations of the Arimoto MI are all equivalent under certain conditions on initial distributions. Similarly to [26, Lemma 7 and 8], it can be shown that the Algorithm S1 achieves a greater increase in the objective function per update than Algorithm S3. Thus, we numerically compared the AO algorithms S1, JO, C, and LP for $\alpha\in(1,\infty)$.
Define a channel
$p_{Y\mid X}$ $:=\begin{bmatrix}0.259&0.463&0.278\\ 0.328&0.172&0.500\\ 0.425&0.225&0.350\end{bmatrix},$ (61)
where $(i,j)$-element of the channel matrix [^2] The elements of the channel matrix $p_{Y\mid X}$ are randomly generated from a uniform distribution. Additional numerical examples were computed using the AO algorithms for other randomly generated channel matrices that did not appear in this paper, wherein the behavior of the algorithms is generally the same. corresponds to the conditional probability $p_{Y\mid X}(j|i),i\in\mathcal{X}=\{1,2,3\},j\in\mathcal{Y}=\{1,2,3\}$. Table II shows the number of iterations $N$ and approximate values of $\alpha$-capacity of the channel $p_{Y\mid X}$, denoted as $F^{(N)}$, for $\alpha\in\{1.03,1.5,2.0,5.0\}$ computed by each algorithm with specific initial conditions. Here, $N$ represents the number of iterations needed to satisfy the stopping conditions. The distributions $u_{X}$ and $u_{X,Y}$ are uniform distributions on $\mathcal{X}$ and $\mathcal{X}\times\mathcal{Y}$, respectively. Figure 1 illustrates transitions of the approximate values of $\alpha$-capacity $F^{(k)}$ as $k$ increases, computed by each algorithm with specific initial conditions. Table II and Figure 1 confirm that Algorithm S1 has the fastest convergence speed.
TABLE II: Approximate values of $\alpha$-capacity and the number of iterations $(F^{(N)},N)$
Algorithm and initial distribution $\alpha=1.03$ $\alpha=1.5$ $\alpha=2.0$ $\alpha=5.0$
Algorithm S1 with $p_{X}^{(0)}=u_{X}$ $(0.054204678,\textbf{851})$ $(0.07617995,\textbf{738})$ $(0.097030615,\textbf{790})$ $(0.183426237,\textbf{20})$
Algorithm JO with $q_{X,Y}^{(0)}=u_{X,Y}$ $(0.054201694,19076)$ $(0.07617965,2744)$ $(0.097030139,2258)$ $(0.183426230,113)$
Algorithm JO with $q_{X,Y}^{(0)}=u_{X}p_{Y\mid X}$ $(0.054201694,19075)$ $(0.07617965,2743)$ $(0.097030139,2257)$ $(0.183426230,112)$
Algorithm C with $p_{X}^{(0)}\tilde{q}_{Y\mid X}^{(0)}=u_{X,Y}$ $(0.054204678,912)$ $(0.07617991,1059)$ $(0.097030445,1347)$ $(0.183426230,97)$
Algorithm C with $p_{X}^{(0)}\tilde{q}_{Y\mid X}^{(0)}=u_{X}p_{Y\mid X}$ $(0.054204678,874)$ $(0.07617991,1055)$ $(0.097030445,1344)$ $(0.183426231,95)$
Algorithm LP with $\tilde{q}_{X,Y}^{(0)}=u_{X,Y}$ $(0.054204763,19540)$ $(0.07618003,3512)$ $(0.097030811,3044)$ $(0.183426237,177)$
Algorithm LP with $\tilde{q}_{X,Y}^{(0)}=u_{X}p_{Y\mid X}$ $(0.054204763,19539)$ $(0.07618003,3511)$ $(0.097030811,3043)$ $(0.183426237,176)$
Refer to caption Refer to caption
(a) $\alpha=1.03$ (b) $\alpha=1.5$
Refer to caption Refer to caption
(c) $\alpha=2.0$ (d) $\alpha=5.0$
Figure 1: Transitions of the approximate value of $\alpha$-capacity $F^{(k)}$ as $k$ increases for (a) $\alpha=1.03$, (b) $\alpha=1.5$, (c) $\alpha=2.0$, and (d) $\alpha=5.0$. The blue solid curve represents Algorithm S1, the orange dashed curve denotes Algorithm JO (modified Jitsumasu–Oohama algorithm), the green dash-dotted curve represents Algorithm C, and the red dash-dot curve indicates Algorithm LP.
VI Conclusion
This paper provides variational characterizations of $\alpha$-MI. Based on these characterizations, several novel AO algorithms for computing $\alpha$-MI and $\alpha$-capacity are presented. The numerical results show that the AO algorithm based on the Sibson MI’s characterization has the fastest convergence speed. Future work will focus on proving the nonasymptotic convergence analysis of these AO algorithms.
Appendix A Proof of Theorem 1
The characterizations in (16) and (17) were originally proved by Arimoto[2, Thm 4] and Kamatsuka et al. [17, Thm 1], respectively. Here, we present a novel proof for (16) using the variational characterizations of the Shannon MI in Proposition 1.
We define the distribution as follows:
$\hat{q}_{X,Y}(x,y)$ $:=\frac{p_{X}(x)^{\frac{1}{\alpha}}p_{Y\mid X}(y\mid x)r_{X\mid Y}(x\mid y)^{1-\frac{1}{\alpha}}}{\sum_{x,y}p_{X}(x)^{\frac{1}{\alpha}}p_{Y\mid X}(y\mid x)r_{X\mid Y}(x\mid y)^{1-\frac{1}{\alpha}}}.$ (62)
Then, for $\alpha\in(0,1)$,
$I_{\alpha}^{\text{S}}(X;Y)\overset{(a)}{=}\min_{\tilde{q}_{X,Y}}\Bigl{\{}\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||\tilde{q}_{X}p_{Y\mid X})+I(\tilde{q}_{X},\tilde{q}_{Y\mid X})$
$\qquad\qquad\qquad\qquad+\frac{1}{1-\alpha}D(\tilde{q}_{X}||p_{X})\Bigr{\}}$ (63)
=$\overset{(b)}{=}\min_{\tilde{q}_{X,Y}}\max_{r_{X\mid Y}}\Bigl{\{}\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||\tilde{q}_{X}p_{Y\mid X})$
$\qquad+\mathbb{E}_{X,Y}^{\tilde{q}_{X,Y}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{\tilde{q}_{X}(X)}\right]+\frac{1}{1-\alpha}D(\tilde{q}_{X}||p_{X})\Bigr{\}}$ (64)
=$\overset{(c)}{=}\max_{r_{X\mid Y}}\min_{\tilde{q}_{X,Y}}\Bigl{\{}\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||\tilde{q}_{X}p_{Y\mid X})$
$\qquad+\mathbb{E}_{X,Y}^{\tilde{q}_{X,Y}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{\tilde{q}_{X}(X)}\right]+\frac{1}{1-\alpha}D(\tilde{q}_{X}||p_{X})\Bigr{\}}$ (65)
$=\max_{r_{X\mid Y}}\min_{\tilde{q}_{X,Y}}\Bigl{\{}D(\tilde{q}_{X,Y}||\hat{q}_{X,Y})+F_{\alpha}^{\text{S1}}(p_{X},r_{X\mid Y})\Bigr{\}}$ (66)
=$\overset{(d)}{=}\max_{r_{X\mid Y}}F_{\alpha}^{\text{S1}}(p_{X},r_{X\mid Y}),$ (67)
where $(a)$ follows from (12), $(b)$ follows from (3), $(c)$ follows from the minimax theorem (see, e.g., [32, Thm 4.2]) [^3] Notably, $\tilde{F}_{\alpha}^{\text{S3}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y}):=\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||\tilde{q}_{X}p_{Y\mid X})+\mathbb{E}_{X,Y}^{\tilde{q}_{X,Y}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{\tilde{q}_{X}(X)}\right]+\frac{1}{1-\alpha}D(\tilde{q}_{X}||p_{X}),\alpha\in(0,1)$ is concave for $r_{X\mid Y}$ (see, e.g. [28, Sec. 10.3.2]) and convex for $\tilde{q}_{X,Y}$, which can be easily verified., and $(d)$ follows from (62) and the properties of the Kullback–Leibler divergence (see, e.g., [33, Thm 2.6.3]). For $\alpha\in(1,\infty)$, Theorem 1 can be shown by replacing $\min_{\tilde{q}_{X,Y}}$ with $\max_{\tilde{q}_{X,Y}}$ in (63)–(67).
Appendix B Proof of Theorem 2
We define the distributions as follows.
$\hat{q}_{Y\mid X}(y\mid x):=\frac{p_{Y\mid X}(y\mid x)r_{X\mid Y}(x\mid y)^{1-\frac{1}{\alpha}}}{\sum_{y}p_{Y\mid X}(y\mid x)r_{X\mid Y}(x\mid y)^{1-\frac{1}{\alpha}}}$ (68)
$\hat{q}_{X}(x):=\frac{p_{X}(x)^{\frac{\alpha}{2\alpha-1}}\left(\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}\right)^{\frac{\alpha}{2\alpha-1}}}{\sum_{x}p_{X}(x)^{\frac{\alpha}{2\alpha-1}}\left(\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}\right)^{\frac{\alpha}{2\alpha-1}}},$ (69)
Then, for $\alpha\in(0,1)$,
$I_{\alpha}^{\text{C}}(X;Y)$
=$\overset{(a)}{=}\min_{\tilde{q}_{Y\mid X}}\left\{I(p_{X},\tilde{q}_{Y\mid X})+\frac{\alpha}{1-\alpha}D(p_{X}\tilde{q}_{Y\mid X}||p_{X}p_{Y\mid X})\right\}$ (70)
=$\overset{(b)}{=}\min_{\tilde{q}_{Y\mid X}}\max_{r_{X\mid Y}}\Biggl{\{}\mathbb{E}_{X,Y}^{p_{X}\tilde{q}_{Y\mid X}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{p_{X}(X)}\right]$
$\qquad\qquad\qquad\qquad+\frac{\alpha}{1-\alpha}D(p_{X}\tilde{q}_{Y\mid X}||p_{X}p_{Y\mid X})\Biggr{\}}$ (71)
=$\overset{(c)}{=}\max_{r_{X\mid Y}}\min_{\tilde{q}_{Y\mid X}}\Biggl{\{}\mathbb{E}_{X,Y}^{p_{X}\tilde{q}_{Y\mid X}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{p_{X}(X)}\right]$
$\qquad\qquad\qquad\qquad+\frac{\alpha}{1-\alpha}D(p_{X}\tilde{q}_{Y\mid X}||p_{X}p_{Y\mid X})\Biggr{\}}$ (72)
$=\max_{r_{X\mid Y}}\min_{\tilde{q}_{Y\mid X}}\Biggl{\{}H(p_{X})$
$+\frac{\alpha}{1-\alpha}\mathbb{E}_{X}^{p_{X}}\left[D\left(\tilde{q}_{Y\mid X}(\cdot\mid X)||\hat{q}_{Y\mid X}(\cdot\mid X)\right)\right]$
$+\frac{\alpha}{\alpha-1}\sum_{x}p_{X}(x)\log\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}\Biggr{\}}$ (73)
[…]
where $\hat{q}_{Y\mid X}(\cdot|x)$ is defined in (68) and $(a)$ follows from [23, Thm 1], $(b)$ follows from (3), $(c)$ follows from the minimax theorem [^4] Notably, $\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X},r_{X\mid Y}):=\mathbb{E}_{X,Y}^{p_{X}\tilde{q}_{Y\mid X}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{p_{X}(X)}\right]+\frac{\alpha}{1-\alpha}D(p_{X}\tilde{q}_{Y\mid X}||p_{X}p_{Y\mid X}),\alpha\in(0,1)$ is concave for $r_{X\mid Y}$ and linear (hence, convex) for $\tilde{q}_{Y\mid X}$. , and $(d)$ follows from the properties of the Kullback–Leibler divergence (see, e.g., [33, Thm 2.6.3]). Eq. (20) for $\alpha\in(0,1)$ can be obtained from (1) and (70) by replacing $I(p_{X},\tilde{q}_{Y\mid X})$ with $\min_{q_{Y}}D(p_{X}\tilde{q}_{Y\mid X}||p_{X}q_{Y})$, which was originally proved by Karakos et al. [18]. For $\alpha\in(1,\infty)$, Theorem 2 can be shown by replacing $\min_{\tilde{q}_{Y\mid X}}$ with $\max_{\tilde{q}_{Y\mid X}}$ in Eqs. (70)–(73).
Appendix C Proof of Theorem 3
We define the distributions as follows.
$\hat{q}_{Y\mid X}(x|y):=\frac{p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}}{\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}},$ (75)
$\hat{q}_{X}(x):=\frac{p_{X}(x)^{\frac{\alpha}{2\alpha-1}}\left(\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}\right)^{\frac{\alpha}{2\alpha-1}}}{\sum_{x}p_{X}(x)^{\frac{\alpha}{2\alpha-1}}\left(\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}\right)^{\frac{\alpha}{2\alpha-1}}}.$ (76)
Then, for $\alpha\in(1,\infty)$,
$I_{\alpha}^{\text{LP}}(X;Y)\overset{(a)}{=}\max_{\tilde{q}_{X,Y}}\left\{\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||p_{X}p_{Y\mid X})+I(\tilde{q}_{X},\tilde{q}_{Y\mid X})\right\}$ (77)
=$\overset{(b)}{=}\max_{\tilde{q}_{X,Y}}\max_{r_{X\mid Y}}\Bigl{\{}\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||p_{X}p_{Y\mid X})$
$\qquad\qquad\qquad+\mathbb{E}_{X,Y}^{\tilde{q}_{X,Y}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{\tilde{q}_{X}(X)}\right]\Bigr{\}}$ (78)
$=\max_{\tilde{q}_{X,Y}}\max_{r_{X\mid Y}}\Bigl{\{}\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}||\tilde{q}_{X}p_{Y\mid X})$
$\qquad+\mathbb{E}_{X,Y}^{\tilde{q}_{X,Y}}\left[\log\frac{r_{X\mid Y}(X\mid Y)}{\tilde{q}_{X}(X)}\right]+D(\tilde{q}_{X}||p_{X})\Bigr{\}}$ (79)
$=\max_{r_{X\mid Y}}\max_{\tilde{q}_{X}}\max_{\tilde{q}_{Y\mid X}}\Bigl{\{}\frac{\alpha}{1-\alpha}\mathbb{E}_{X}^{\tilde{q}_{X}}\left[D(\tilde{q}_{Y\mid X}(\cdot\mid X)||\hat{q}_{Y\mid X}(\cdot\mid X))\right]$
$+H(\tilde{q}_{Y})+\frac{\alpha}{1-\alpha}D(\tilde{q}_{X}||p_{X})\Bigr{\}}$ (80)
[…]
where $(a)$ follows from (13), $(b)$ follows from (3), and $(c)$ and $(d)$ follow from the properties of the Kullback–Leibler divergence. For $\alpha\in(0,1)$, Eq. (24) for $\alpha\in(0,1)$ can be obtained by replacing $\max_{\tilde{q}_{X,Y}}$ with $\min_{\tilde{q}_{X,Y}}$ and $I(\tilde{q}_{X},\tilde{q}_{Y\mid X})$ with $\min_{q_{Y}}D(\tilde{q}_{X}\tilde{q}_{Y\mid X}||\tilde{q}_{X}q_{Y})$ in (77).
Appendix D Proof of Proposition 4
Eqs. (32), (36), and (37) immediately follow from the proof of Theorem 3 and Proposition 1. Here, we prove only (33). Let $\bar{q}_{X}$ and $\bar{q}_{Y\mid X}$ be distributions defined in (34) and (35), respectively. Let $\tilde{q}_{X}(x):=\sum_{y}\tilde{q}_{X,Y}(x,y)$. Then,
$F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},q_{Y})=\frac{1}{1-\alpha}\mathbb{E}_{X,Y}^{\tilde{q}_{X}}\left[D(\tilde{q}_{Y\mid X}(\cdot\mid X)||\bar{q}_{Y\mid X}(\cdot\mid X))\right]$
$+\frac{1}{\alpha-1}\sum_{x}\tilde{q}_{X}(x)\log\sum_{y}p_{Y\mid X}(y\mid x)^{\alpha}q_{Y}(y)^{1-\alpha}$
$+\frac{\alpha}{1-\alpha}D(\tilde{q}_{X}||p_{X})$ (83)
$\geq\frac{1}{\alpha-1}\sum_{x}\tilde{q}_{X}(x)\log\sum_{y}p_{Y\mid X}(y\mid x)^{\alpha}q_{Y}(y)^{1-\alpha}$
$+\frac{\alpha}{1-\alpha}D(\tilde{q}_{X}||p_{X})$ (84)
$=\frac{\alpha}{1-\alpha}D(\tilde{q}_{X}||\bar{q}_{X})$
$+\frac{\alpha}{\alpha-1}\log\sum_{x}p_{X}(x)\left(\sum_{y}p_{Y\mid X}(y\mid x)^{\alpha}q_{Y}(y)^{1-\alpha}\right)^{\frac{1}{\alpha}}$ (85)
$\geq\frac{\alpha}{\alpha-1}\log\sum_{x}p_{X}(x)\left(\sum_{y}p_{Y\mid X}(y\mid x)^{\alpha}q_{Y}(y)^{1-\alpha}\right)^{\frac{1}{\alpha}},$ (86)
where the equality in (84) is achieved when $\tilde{q}_{Y\mid X}=\bar{q}_{Y\mid X}$; the equality in (86) is achieved when $\tilde{q}_{X}=\bar{q}_{X}$.
Appendix E Proof of Lemma 2
Proof.
Let $\tilde{q}_{X,Y}^{\text{opt}}$ and $p_{X}^{\text{opt}}$ be the optimal distributions that attain the Lapidoth–Pfister capacity $C_{\alpha}^{\text{LP}}=\max_{p_{X}}\max_{\tilde{q}_{X,Y}}\max_{r_{X\mid Y}}\tilde{F}_{\alpha}^{\text{LP}}(p_{X},\tilde{q}_{X,Y},r_{X\mid Y})$. Let $\tilde{q}_{X,Y}^{(k)}(x,y)=\tilde{q}_{X}^{(k)}(x)\tilde{q}_{Y\mid X}^{(k)}(y|x),\tilde{q}_{Y}^{(k)}(y):=\sum_{x}\tilde{q}_{X,Y}^{(k)}(x,y)$ and $\tilde{q}_{X,Y}^{\text{opt}}(x,y)=\tilde{q}_{X}^{\text{opt}}(x)\tilde{q}_{Y\mid X}^{\text{opt}}(y|x),\tilde{q}_{Y}^{\text{opt}}(y):=\sum_{x}\tilde{q}_{X,Y}^{\text{opt}}(x,y)$. Then, from Theorem 3 and Proposition 4, we can obtain the following:
$\tilde{F}_{\alpha}^{\text{LP}}(p_{X}^{(k+1)},\tilde{q}_{X,Y}^{(k+1)},r_{X\mid Y}^{(k)})=\max_{\tilde{q}_{X,Y}}\tilde{F}_{\alpha}^{\text{LP}}(p_{X}^{(k+1)},\tilde{q}_{X,Y},r_{X\mid Y}^{(k)})$ (87)
=$\overset{(a)}{=}\frac{2\alpha-1}{\alpha-1}$
×$\times\log\sum_{x}p_{X}^{(k+1)}(x)^{\frac{\alpha}{2\alpha-1}}\left(\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}(x|y)^{1-\frac{1}{\alpha}}\right)^{\frac{\alpha}{2\alpha-1}},$ (88)
$\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}^{(k)}(x|y)^{1-\frac{1}{\alpha}}\overset{(b)}{=}\frac{p_{Y\mid X}(y|x)r_{X\mid Y}^{(k)}(x|y)^{1-\frac{1}{\alpha}}}{\tilde{q}_{Y\mid X}^{(k+1)}(y|x)}$ (89)
=$\overset{(c)}{=}\frac{p_{Y\mid X}(y|x)\tilde{q}_{X,Y}^{(k)}(x,y)^{1-\frac{1}{\alpha}}}{\tilde{q}_{Y\mid X}^{(k+1)}(y|x)}\cdot\frac{1}{\tilde{q}_{Y}^{(k)}(y)^{1-\frac{1}{\alpha}}},$ (90)
$\sum_{x}p_{X}^{(k+1)}(x)^{\frac{\alpha}{2\alpha-1}}\left(\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}^{(k)}(x|y)^{1-\frac{1}{\alpha}}\right)^{\frac{\alpha}{2\alpha-1}}$
=$\overset{(d)}{=}\frac{p_{X}^{(k+1)}(x)^{\frac{\alpha}{2\alpha-1}}\left(\sum_{y}p_{Y\mid X}(y|x)r_{X\mid Y}^{(k)}(x|y)^{1-\frac{1}{\alpha}}\right)^{\frac{\alpha}{2\alpha-1}}}{\tilde{q}_{X}^{(k+1)}(x)}$ (91)
=$\overset{(e)}{=}\frac{p_{X}^{(k+1)}(x)^{\frac{\alpha}{2\alpha-1}}}{\tilde{q}_{X}^{(k+1)}(x)}$
×$\times\left(\frac{p_{Y\mid X}(y|x)\tilde{q}_{X,Y}^{(k)}(x,y)^{1-\frac{1}{\alpha}}}{\tilde{q}_{Y\mid X}^{(k+1)}(y|X)}\cdot\frac{1}{\tilde{q}_{Y}^{(k)}(y)^{1-\frac{1}{\alpha}}}\right)^{\frac{\alpha}{2\alpha-1}}$ (92)
$=\frac{p_{X}^{(k+1)}(x)^{\frac{\alpha}{2\alpha-1}}p_{Y\mid X}(y|x)^{\frac{\alpha}{2\alpha-1}}\tilde{q}_{X,Y}^{(k)}(x,y)^{\frac{\alpha-1}{2\alpha-1}}}{\tilde{q}_{X}^{(k+1)}(x)\tilde{q}_{Y\mid X}^{(k+1)}(y|x)^{\frac{\alpha}{2\alpha-1}}\tilde{q}_{Y}^{(k)}(y)^{\frac{\alpha-1}{2\alpha-1}}}$ (93)
for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$, where $(a)$ follows from (27), $(b)$ follows from (37) and (39), $(c)$ follows from (36), $(d)$ follows from (37) and (38), and $(e)$ follows from (90). We can substitute (93) into (88) to obtain
$F_{\alpha}^{\text{LP}}(p_{X}^{(k+1)},\tilde{q}_{X,Y}^{(k+1)},r_{X\mid Y}^{(k)})$
$=\frac{2\alpha-1}{\alpha-1}\log\frac{p_{X}^{(k+1)}(x)^{\frac{\alpha}{2\alpha-1}}p_{Y\mid X}(y|x)^{\frac{\alpha}{2\alpha-1}}\tilde{q}_{X,Y}^{(k)}(x,y)^{\frac{\alpha-1}{2\alpha-1}}}{\tilde{q}_{X}^{(k+1)}(x)\tilde{q}_{Y\mid X}^{(k+1)}(y|x)^{\frac{\alpha}{2\alpha-1}}\tilde{q}_{Y}^{(k)}(y)^{\frac{\alpha-1}{2\alpha-1}}}$ (94)
for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Then, we can use (94) to obtain the upper bound of the approximation error as follows:
$C_{\alpha}^{\text{LP}}-\tilde{F}_{\alpha}^{\text{LP}}(p_{X}^{(k+1)},\tilde{q}_{X,Y}^{(k+1)},r_{X\mid Y}^{(k)})$
$=\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}^{\text{opt}}||p_{X}^{\text{opt}}p_{Y\mid X})+I(\tilde{q}_{X}^{\text{opt}},\tilde{q}_{Y\mid X}^{\text{opt}})$
$-\frac{2\alpha-1}{\alpha-1}\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)F_{\alpha}^{\text{LP}}(p_{X}^{(k+1)},\tilde{q}_{X,Y}^{(k+1)},r_{X\mid Y}^{(k)})$ (95)
$=\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}^{\text{opt}}||p_{X}^{\text{opt}}p_{Y\mid X})+I(\tilde{q}_{X}^{\text{opt}},\tilde{q}_{Y\mid X}^{\text{opt}})-\frac{2\alpha-1}{\alpha-1}$
×$\times\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)\log\frac{p_{X}^{(k+1)}(x)^{\frac{\alpha}{2\alpha-1}}p_{Y\mid X}(y|x)^{\frac{\alpha}{2\alpha-1}}\tilde{q}_{X,Y}^{(k)}(x,y)^{\frac{\alpha-1}{2\alpha-1}}}{\tilde{q}_{X}^{(k+1)}(x)\tilde{q}_{Y\mid X}^{(k+1)}(y|x)^{\frac{\alpha}{2\alpha-1}}\tilde{q}_{Y}^{(k)}(y)^{\frac{\alpha-1}{2\alpha-1}}}$ (96)
=$\overset{(f)}{=}\frac{\alpha}{1-\alpha}\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)\log\frac{\tilde{q}_{Y\mid X}^{\text{opt}}(y|x)^{\frac{1}{\alpha}}}{\tilde{q}_{Y\mid X}^{(k)}(y|x)^{\frac{1}{\alpha}}}\cdot\frac{\tilde{q}_{X}^{(k)}(x)\tilde{q}_{Y\mid X}^{(k)}(y|x)}{\tilde{q}_{X}^{(k+1)}(x)\tilde{q}_{Y\mid X}^{(k+1)}(y|x)}$
$\cdot\frac{\tilde{q}_{X}^{(k+1)}(x)^{\frac{1}{\alpha}}-1}{\tilde{q}_{X}^{(k)}(x)^{\frac{1}{\alpha}}-1}\cdot\frac{\tilde{q}_{Y}^{(k)}(y)^{\frac{1}{\alpha}-1}}{\tilde{q}_{Y}^{\text{opt}}(y)^{\frac{1}{\alpha}-1}}$ (97)
$=\frac{1}{1-\alpha}\mathbb{E}_{X}^{\tilde{q}_{X}^{\text{opt}}}\left[D(\tilde{q}_{Y\mid X}^{\text{opt}}(\cdot\mid X)||\tilde{q}_{Y\mid X}(\cdot\mid X))\right]$
$+\frac{\alpha}{\alpha-1}\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}\log\frac{\tilde{q}_{X,Y}^{k+1}(x,y)}{\tilde{q}_{X,Y}^{(k)}(x,y)}$
$+\sum_{x}\tilde{q}_{X}^{\text{opt}}(x)\log\frac{\tilde{q}_{X}^{(k+1)}(x)}{\tilde{q}_{X}^{(k)}(x)}-D(\tilde{q}_{Y}^{\text{opt}}||\tilde{q}_{Y}^{(k)})$ (98)
$\overset{(g)}{\leq}\frac{\alpha}{\alpha-1}\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}\log\frac{\tilde{q}_{X,Y}^{k+1}(x,y)}{\tilde{q}_{X,Y}^{(k)}(x,y)}+\sum_{x}\tilde{q}_{X}^{\text{opt}}(x)\log\frac{\tilde{q}_{X}^{(k+1)}(x)}{\tilde{q}_{X}^{(k)}(x)},$ (99)
where $(f)$ follows from (44), i.e., $p_{X}^{\text{opt}}=\tilde{q}_{X}^{\text{opt}}$ and $p_{X}^{(k)}=\tilde{q}_{X}^{(k)}$, and $(g)$ follows from the nonnegativity of the Kullback–Leibler divergence and $\frac{1}{1-\alpha}<0$ for $\alpha\in(1,\infty)$
Appendix F Proof of Lemma 4
Proof.
For convenience, let $\tilde{\gamma}^{(k)}(y):=\sum_{x}p_{X}(x)\tilde{q}_{Y\mid X}^{(k)}(y|x)$ and $\tilde{\gamma}^{\text{opt}}(y):=\sum_{x}p_{X}(x)\tilde{q}_{Y\mid X}^{\text{opt}}(y|x)$. Then, Theorem 2 and Proposition 3 can be used to obtain the following:
$\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X}^{(k+1)},r_{X\mid Y}^{(k)})=\max_{\tilde{q}_{Y\mid X}}\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X},r_{X\mid Y}^{(k)})$ (100)
=$\overset{(a)}{=}H(p_{X})$
$+\frac{\alpha}{\alpha-1}\sum_{x}p_{X}(x)\log\sum_{y}p_{Y\mid X}(y\mid x)r_{X\mid Y}^{(k)}(x\mid y)^{1-\frac{1}{\alpha}},$ (101)
$\sum_{y}p_{Y\mid X}(y\mid x)r_{X\mid Y}^{(k)}(x\mid y)^{1-\frac{1}{\alpha}}$
=$\overset{(b)}{=}\frac{p_{Y\mid X}(y\mid x)r_{X\mid Y}^{(k)}(x\mid y)^{1-\frac{1}{\alpha}}}{\tilde{q}_{Y\mid X}^{(k+1)}(y\mid x)}$ (102)
=$\overset{(c)}{=}\frac{p_{Y\mid X}(y\mid x)\left(\frac{p_{X}(x)\tilde{q}_{Y\mid X}^{(k)}(y\mid x)}{\tilde{\gamma}^{(k)}(y)}\right)^{1-\frac{1}{\alpha}}}{\tilde{q}_{Y\mid X}^{(k+1)}(y\mid x)}$ (103)
$=\frac{p_{Y\mid X}(y\mid x)p_{X}(x)^{1-\frac{1}{\alpha}}\tilde{q}_{Y\mid X}^{(k)}(y\mid x)^{1-\frac{1}{\alpha}}}{\tilde{q}_{Y\mid X}^{(k+1)}(y\mid x)}\times\frac{1}{\tilde{\gamma}^{(k)}(y)^{1-\frac{1}{\alpha}}}.$ (104)
for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$. We can substitute (104) into (101) to obtain:
$\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X}^{(k+1)},r_{X\mid Y}^{(k)})=H(p_{X})+\frac{\alpha}{\alpha-1}\sum_{x}p_{X}(x)$
$\log\frac{p_{Y\mid X}(y\mid x)p_{X}(x)^{1-\frac{1}{\alpha}}\tilde{q}_{Y\mid X}^{(k)}(y\mid x)^{1-\frac{1}{\alpha}}}{\tilde{q}_{Y\mid X}^{(k+1)}(y\mid x)}\cdot\frac{1}{\tilde{\gamma}^{(k)}(y)^{1-\frac{1}{\alpha}}}$ (105)
$=\frac{\alpha}{\alpha-1}\sum_{x}p_{X}(x)\log\frac{\tilde{q}_{Y\mid X}^{(k)}(y\mid x)^{1-\frac{1}{\alpha}}}{\tilde{q}_{Y\mid X}^{(k+1)}(y\mid x)}\cdot\frac{p_{Y\mid X}(y\mid x)}{\tilde{\gamma}^{(k)}(y)^{1-\frac{1}{\alpha}}}$ (106)
for all $y\in\mathcal{Y}$. Then, we can use (106) to obtain the upper bound of the approximation error as follows:
$I_{\alpha}^{\text{C}}(X;Y)-\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X}^{(k+1)},r_{X\mid Y}^{(k)})$
$=I(p_{X},\tilde{q}_{Y\mid X}^{\text{opt}})+\frac{\alpha}{1-\alpha}D(p_{X}\tilde{q}_{Y\mid X}^{\text{opt}}||p_{X}p_{Y\mid X})$
$\qquad\qquad-\sum_{y}\tilde{q}_{Y\mid X}^{\text{opt}}(y\mid x)\tilde{F}_{\alpha}^{\text{C}}(\tilde{q}_{Y\mid X}^{(k+1)},r_{X\mid Y}^{(k)})$ (107)
$=\frac{\alpha}{1-\alpha}D(p_{X}\tilde{q}_{Y\mid X}^{\text{opt}}||p_{X}p_{Y\mid X})+I(p_{X},\tilde{q}_{Y\mid X}^{\text{opt}})-\frac{\alpha}{\alpha-1}$
×$\times\sum_{x,y}p_{X}(x)\tilde{q}_{Y\mid X}^{\text{opt}}(y\mid x)\log\frac{\tilde{q}_{Y\mid X}^{(k)}(y\mid x)^{1-\frac{1}{\alpha}}}{\tilde{q}_{Y\mid X}^{(k+1)}(y\mid x)}\cdot\frac{p_{Y\mid X}(y\mid x)}{\tilde{\gamma}^{(k)}(y)^{1-\frac{1}{\alpha}}}$ (108)
$=\frac{\alpha}{\alpha-1}\sum_{x,y}p_{X}(x)\tilde{q}_{Y\mid X}^{\text{opt}}(y\mid x)\log\frac{\tilde{q}_{Y\mid X}^{(k+1)}(y\mid x)}{\tilde{q}_{Y\mid X}^{(k)}(y\mid x)}$
$+\frac{1}{1-\alpha}\mathbb{E}_{X}^{p_{X}}(x)\left[D(\tilde{q}_{Y\mid X}^{\text{opt}}(\cdot\mid X)||\tilde{q}_{Y\mid X}^{(k)}(\cdot\mid X))\right]-D(\tilde{\gamma}^{\text{opt}}||\tilde{\gamma}^{(k)})$ (109)
$\overset{(d)}{\leq}\frac{\alpha}{\alpha-1}\sum_{x,y}p_{X}(x)\tilde{q}_{Y\mid X}^{\text{opt}}(y\mid x)\log\frac{\tilde{q}_{Y\mid X}^{(k+1)}(y\mid x)}{\tilde{q}_{Y\mid X}^{(k)}(y\mid x)},$ (110)
where $(d)$ follows from the nonnegativity of the Kullback–Leibler divergence and $\frac{1}{1-\alpha}<0$ for $\alpha\in(1,\infty)$. ∎
Appendix G Proof of Lemma 6
Proof.
Let $\tilde{q}_{X,Y}^{(k)}=\tilde{q}_{X}^{(k)}\tilde{q}_{Y\mid X}^{(k)}$ and $\tilde{q}_{Y}^{(k)}(y):=\sum_{x}\tilde{q}_{X,Y}^{(k)}(x,y)$. Then, from Theorem 3 and Proposition 4, we can obtain the following:
$F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y}^{(k+1)},q_{Y}^{(k)})=\max_{\tilde{q}_{X,Y}}F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y},q_{Y}^{(k)})$
=$\overset{(a)}{=}\frac{\alpha}{\alpha-1}\log\sum_{x}p_{X}(x)\left(\sum_{y}p_{Y\mid X}(y|x)^{\alpha}q_{Y}^{(k)}(y)^{1-\alpha}\right)^{\frac{1}{\alpha}},$ (111)
$\sum_{y}p_{Y\mid X}(y|x)^{\alpha}q_{Y}^{k}(y)^{1-\alpha}\overset{(b)}{=}\frac{p_{Y\mid X}(y|x)^{\alpha}q_{Y}^{(k)}(y)^{1-\alpha}}{\tilde{q}_{Y\mid X}^{(k)}(y|x)}$ (112)
=$\overset{(c)}{=}\frac{p_{Y\mid X}(y|x)^{\alpha}\tilde{q}_{Y}^{(k)}(y)^{1-\alpha}}{\tilde{q}_{Y\mid X}^{(k)}(y|x)},$ (113)
$\sum_{x}p_{X}(x)\left(\sum_{y}p_{Y\mid X}(y|x)^{\alpha}q_{Y}^{(k)}(y)^{1-\alpha}\right)^{\frac{1}{\alpha}}$
=$\overset{(d)}{=}\frac{p_{X}(x)\left(\sum_{y}p_{Y\mid X}(y|x)^{\alpha}q_{Y}^{(k)}(y)^{1-\alpha}\right)^{\frac{1}{\alpha}}}{\tilde{q}_{X}^{(k+1)}(x)}$ (114)
=$\overset{(e)}{=}\frac{p_{X}(x)\left(\frac{p_{Y\mid X}(y|x)^{\alpha}\tilde{q}_{Y}^{(k)}(y)^{1-\alpha}}{\tilde{q}_{Y\mid X}^{(k)}(y|x)}\right)^{\frac{1}{\alpha}}}{\tilde{q}_{X}^{(k+1)}(x)}$ (115)
$=\frac{p_{X}(x)p_{Y\mid X}(y|x)\tilde{q}_{Y}^{(k)}(y)^{\frac{1}{\alpha}-1}}{\tilde{q}_{X}^{(k+1)}(x)\tilde{q}_{Y\mid X}^{(k)}(y|x)^{\frac{1}{\alpha}}},$ (116)
where $(a)$ follows from Theorem 3, Proposition 4, and a simple calculation, $(b)$ follows from (33) and (35); $(c)$ follows from (32); $(d)$ follows from (33) and (34); and $(e)$ follows from (113).
We can substitute (116) into (111) to obtain
$F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y}^{(k+1)},q_{Y}^{(k)})=\frac{\alpha}{\alpha-1}\log\frac{p_{X}(x)p_{Y\mid X}(y|x)\tilde{q}_{Y}^{(k)}(y)^{\frac{1}{\alpha}-1}}{\tilde{q}_{X}^{(k+1)}(x)\tilde{q}_{Y\mid X}^{(k)}(y|x)^{\frac{1}{\alpha}}}$ (117)
for all $(x,y)\in\mathcal{X}\times\mathcal{Y}$. Then, we can use (117) to obtain the upper bound of the approximation error as follows:
$F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y}^{(k+1)},q_{Y}^{(k)})-I_{\alpha}^{\text{LP}}(X;Y)$
$=\sum_{x,y}\tilde{q}_{X,Y}^{\text{opt}}(x,y)F_{\alpha}^{\text{LP}}(\tilde{q}_{X,Y}^{(k+1)},q_{Y}^{(k)})$
$-\left(\frac{\alpha}{1-\alpha}D(\tilde{q}_{X,Y}^{\text{opt}}||p_{X}p_{Y\mid X})+I(\tilde{q}_{X}^{\text{opt}},\tilde{q}_{Y\mid X}^{\text{opt}})\right)$ (118)
$=D(\tilde{q}_{Y}^{\text{opt}}||\tilde{q}_{Y}^{(k)})-\frac{1}{1-\alpha}D(\tilde{q}_{X,Y}^{\text{opt}}||\tilde{q}_{X}^{\text{opt}}\tilde{q}_{Y\mid X}^{(k)})$
$-\frac{\alpha}{1-\alpha}D(\tilde{q}_{X}^{\text{opt}}||\tilde{q}_{X}^{(k+1)})$ (119)
$=\frac{1-2\alpha}{1-\alpha}D(\tilde{q}_{Y}^{\text{opt}}||\tilde{q}_{Y}^{(k)})-\frac{1}{1-\alpha}D(\tilde{q}_{X,Y}^{\text{opt}}||\tilde{q}_{X}^{\text{opt}}\tilde{q}_{Y\mid X}^{(k)})$
$+\frac{\alpha}{1-\alpha}\left(D(\tilde{q}_{Y}^{\text{opt}}||\tilde{q}_{Y}^{(k)})-D(\tilde{q}_{X}^{\text{opt}}||\tilde{q}_{X}^{(k+1)})\right)$ (120)
$\overset{(f)}{\leq}\frac{\alpha}{1-\alpha}\left(D(\tilde{q}_{Y}^{\text{opt}}||\tilde{q}_{Y}^{(k)})-D(\tilde{q}_{X}^{\text{opt}}||\tilde{q}_{X}^{(k+1)})\right)$ (121)
$\overset{(g)}{\leq}\frac{\alpha}{1-\alpha}\left(D(\tilde{q}_{X}^{\text{opt}}||\tilde{q}_{X}^{(k)})-D(\tilde{q}_{X}^{\text{opt}}||\tilde{q}_{X}^{(k+1)})\right)$ (122)
$=\frac{\alpha}{1-\alpha}\sum_{x}\tilde{q}_{X}^{\text{opt}}(x)\log\frac{\tilde{q}_{X}^{(k+1)}(x)}{\tilde{q}_{X}^{(k)}(x)},$ (123)
where $(f)$ follows from the nonnegativity of the Kullback–Leibler divergence and $\frac{1-2\alpha}{1-\alpha}\leq 0$ for $\alpha\in[1/2,1]$ and $(g)$ follows from the data-processing inequality (also known as monotonicity) of the Kullback–Leibler divergence (see, e.g., [24, Thm 2.17]). ∎